OpenStudy (anonymous):

p=$89700, r=9%, t=2 months[Assume 360 days in a year}

OpenStudy (anonymous):

i have to show work

OpenStudy (paxpolaris):

compounded daily/continuously/....? or simple interest?

OpenStudy (anonymous):

simple intrest

OpenStudy (paxpolaris):

Does the question ask for the interest of the total amount after 2 months?

OpenStudy (paxpolaris):

Start with the formula. which is ...?

OpenStudy (anonymous):

it just say t=2 months

OpenStudy (anonymous):

i=prt

OpenStudy (anonymous):

?

OpenStudy (paxpolaris):

Right. \[\large I=P \cdot r \cdot t\] I is the interest you are trying to calculate. P is the principal (the money you invest / start with) r is the annual interest rate... \(9\%=\Large\frac9{100}\) or 0.09 t is the time in years.. 2months=\(\Large{2 \over 12}\) or \(\Large{1\over 6}\)

OpenStudy (anonymous):

i = p * r * t = 89, 700 * .09 * 1/6

OpenStudy (anonymous):

a=1345.50 b=134550.00 c=16146.00 d=4036.50

OpenStudy (anonymous):

i got it its a thank u

OpenStudy (paxpolaris):

\[\Large I=897\cancel{00}\cdot {\cancel93 \over \cancel{100}} \cdot \frac 1 {\cancel{6}2}\]

OpenStudy (anonymous):

p=$7101, r=1.25, t=11months

OpenStudy (anonymous):

@jordan231 one question per post, pls! And don't forget to click on the blue Best Answer!

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

I mean for PaxPolaris is the person guiding you through the end!